# Equivalence of Privacy and Stability with Generalization Guarantees in Quantum Learning

### v1

@@ IV-A Trusted Setting and One-Neighbor $(\varepsilon,\delta)$-DP
This section examines the connection between privacy and generalization in quantum learning. Building on Section III, which linked information-theoretic stability to generalization, we demonstrate that differential privacy enforces this stability. Using the framework of Figure 2, we introduce the Investigator as the recipient of the output system $B$, generated by a Trusted Data Processor from the Respondent’s raw data ($S,\bm{\mathfrak{Te}},\bm{\mathfrak{Tr}}$). To prevent the reconstruction of individual entries, the Processor ensures the algorithm satisfies differential privacy, requiring output invariance under single-entry modifications. This constraint is mathematically equivalent to algorithmic stability (Definition 1), confirming privacy as a sufficient condition for generalization. We formalize this indistinguishability requirement below.
@@
    The support consistency condition (10) is not strictly necessary for all stability measures. See Remark 4 for more details.
@@
It is important to note that if stability is defined via the trace distance $\|\sigma^{S\bm{\mathfrak{Te}}WB^{\prime}}_{\mathcal{N}}-(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}})\|_{1}$ rather than mutual information, (10) is not required, and similar bounds can be derived using our proof techniques. Note that for $\delta=0$, (10) is automatically satisfied.
@@ IV-B Privacy Implies Stability: A Mutual-Information Bound
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (12)
where, $n$ is the length of the training data and for some constant $m\in(0,1],$ $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\log\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ and has a property that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$.
@@
For $\varepsilon\in\left.[0,\frac{1}{n}\right)$, consider a learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 6 and the condition (11). Then,
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
@@
Furthermore, Theorem 2 strictly generalizes [17, Proposition 2]: by taking a trivial quantum system ($\dim(B^{\prime})=1$) and setting $\delta=0$, the overhead $h_{\left|\mathcal{Z}\right|}$ vanishes, recovering the classical stability bound $(|\mathcal{Z}|-1)\log(ne\varepsilon)$.
@@
The upper-bound obtained in Theorem 2 is independent of $P_{S}$ and thus, Theorem 2 implies that if a quantum learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ satisfies Definition 6, then $\mathcal{N}$ is $\left((|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+\right$. $\left.h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right)$-stable (see Definition 1). A similar observation also follows for Corollaries 3 and 4.
In Appendix B, we present a detailed comparison of Theorem 2 with [15, Proposition 10]. Additionally, in Appendix C, we provide an in-depth comparison of Theorem 2 with the results of [6, Appendix C.7] in the setting of untrusted Data Processors, a topic we will elaborate on in the subsequent section.
@@ IV-C From Stability to Generalization: DP Generalization Guarantees
Consider a quantum learning algorithm $\mathcal{N}$ that is $1$-neighbor $(\varepsilon,\delta)$-DP support consistent (see Definition 6) with $\varepsilon\in[\frac{1}{n},1)$ and satisfies condition (11). If the loss operator satisfies (6) (or equivalently (7)) mentioned in Definition 5, for some $\alpha\in(0,\infty)$, then,
$\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{2\alpha^{2}\mathcal{I}_{bound}},$
where $\mathcal{I}_{bound}=\left[(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right]$ is the upper bound on the mutual information derived in Theorem 2.
@@ V-A Untrusted Data Processor Model
In this setting, the Respondent does not grant the Data Processor access to the raw classical data $s$ directly. Instead, the Respondent provides access only to the set of encoded quantum states $\{\rho_{s}\}_{s}$. The Processor is tasked with running a learning algorithm to produce a hypothesis $w$. Since the Data Processor is untrusted, the learning algorithm is modeled as a single, fixed completely positive, trace-preserving (CP-TP) map $\mathcal{N}$ that must be independent of the specific input index $s$. The total state generated at the output of this process is, $\mathcal{N}(\rho_{s}):=\sum_{w\in\mathcal{W}}\mathcal{N}_{w}(\rho_{s})\otimes|w\rangle\langle w|,$ where each $\mathcal{N}_{w}$ is a completely positive trace non-increasing map summing to $\mathcal{N}$. To ensure privacy, the Respondent mandates that this map $\mathcal{N}$ must satisfy differential privacy constraints with respect to neighboring inputs. The formal definition of such a $1$-neighbor $(\varepsilon,\delta)$-DP support-consistent learning algorithm for an untrusted Data Processor is provided in Definition 8 of Appendix A.
@@ V-B Information-Theoretic Admissibility (ITA): Motivation and Definition
If such a scenario is possible, the privacy guarantees calculated based on $\mathcal{N}$ (as mentioned in Definition 8) are rendered void, as the Data Processor effectively holds the information content of $\mathcal{N}^{\prime}$. To formalize this, we introduce the concept of ordering of informativeness between learning algorithms (Definition 9, Appendix A). We say that an algorithm $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$ with respect to the ensemble $\{\rho_{s}\}_{s}$ if there exists a post-processing CP-TP map $\Gamma$ (a "simulation map") such that,
@@ V-C Quantum Advantage: Privacy under ITA
V-C1 The Collapse of Classical Privacy:
In the classical domain, the encoded states $\rho_{s}$ effectively behave as probability distributions (or commutative states). In such a scenario, the lemma below shows that simultaneous ITA (optimality) and Differential Privacy (indistinguishability) are impossible.
@@ V-C1 The Collapse of Classical Privacy:
See Appendix G
The implication of Lemma 1 is severe: classical ITA algorithms permit full reconstruction of the raw data. Since under ITA the Data Processor effectively holds the raw data, the output-based guarantees of Definition 8 are insufficient. Therefore, for classical ITA algorithms, the privacy condition in Definition 8 must be strengthened to require indistinguishability on the raw states $\rho_{s}$ directly, effectively substituting $\mathcal{N}(\rho_{s})$ with $\rho_{s}$.
V-C2 Quantum Privacy via Non-Commutativity:
In the quantum regime, the raw inputs $\{\rho_{s}\}_{s}$ may be non-commuting. Quantum mechanics dictates that non-orthogonal states cannot be perfectly distinguished, even by the optimal measurement. Therefore, an algorithm can be ITA—extracting the maximum physically permissible information—without yielding enough information to identify $s$ perfectly.
@@ V-C2 Quantum Privacy via Non-Commutativity:
Consider the states $\rho_{z}=|\phi_{z,p}\rangle\langle\phi_{z,p}|$ with $|\phi_{z,p}\rangle:=\sqrt{1-p}$ $|0\rangle+(-1)^{z}\sqrt{p}|1\rangle$. For a dataset $s=(z_{1},\dots,z_{n})$, define the encoded state $|e_{s}\rangle:=\bigotimes_{j=1}^{n}|\phi_{z_{j},1/2}\rangle$.Let $P_{k}$ be the projection onto the subspace spanned by $\{|e_{s}\rangle:|s|_{1}=k\}$ (i.e., the subspace of states corresponds to all strings $s$ with Hamming weight $k$). The untrusted Data Processor applies the projective measurement map $\{\mathcal{N}_{k}\}_{k}$ defined by $\mathcal{N}_{k}(\rho):=P_{k}\rho P_{k}$.

### v2

@@ IV-A Trusted Setting and One-Neighbor $(\varepsilon,\delta)$-DP
This section examines the connection between privacy and generalization in quantum learning. Building on Section III, which linked information-theoretic stability to generalization, we demonstrate that differential privacy enforces this stability. Using the framework of Figure 2, we introduce the Investigator as the recipient of the output system $B$, generated by a Trusted Data Processor from the Respondent’s raw data ($S,\bm{\mathfrak{Te}},\bm{\mathfrak{Tr}}$).
To prevent the reconstruction of individual entries, the Processor ensures the algorithm satisfies differential privacy, requiring output invariance under single-entry modifications. This constraint is mathematically equivalent to algorithmic stability (Definition 1), confirming privacy as a sufficient condition for generalization. We formalize this indistinguishability requirement below.
@@
The support consistency condition (11) is automatically satisfied in the pure differential privacy regime ($\delta=0$).
@@ IV-B Privacy Implies Stability: A Mutual-Information Bound
$I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (13)
where, $n$ is the length of the training data and for some constant $m\in(0,1],$ $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\ln\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ and has a property that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$.
@@
For $\varepsilon\in\left.[0,\frac{1}{n}\right)$, consider a learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$, which satisfies the properties mentioned in Definition 6 and the condition (12). Then, $I[S;WB^{\prime}]\leq(|\mathcal{Z}|-1)\varepsilon n+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
@@
Furthermore, Theorem 4 strictly generalizes [23, Proposition 2]: by taking a trivial quantum system ($\dim(B^{\prime})=1$) and setting $\delta=0$, the overhead $h_{\left|\mathcal{Z}\right|}$ vanishes, recovering the classical stability bound $(|\mathcal{Z}|-1)\ln(ne\varepsilon)$.
@@
The upper-bound obtained in Theorem 4 is independent of $P_{S}$ and thus, Theorem 4 implies that if a quantum learning algorithm $\mathcal{N}=\left\{\mathcal{N}^{(s)}\right\}_{s\in\mathcal{S}}$ satisfies Definition 6, then $\mathcal{N}$ is $\left((|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+\right$. $\left.h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right)$-stable (see Definition 1). A similar observation also follows for Corollaries 3 and 4.
In Appendix F-A, we present a detailed comparison of Theorem 4 with [20, Proposition 10]. Additionally, in Appendix F-B, we provide an in-depth comparison of Theorem 4 with the results of [8, Appendix C.7] in the setting of untrusted Data Processors, a topic we will elaborate on in the subsequent section.
@@ IV-C From Stability to Generalization: DP Generalization Guarantees
Consider a quantum learning algorithm $\mathcal{N}$ that is $1$-neighbor $(\varepsilon,\delta)$-DP support consistent (see Definition 6) with $\varepsilon\in[\frac{1}{n},1)$ and satisfies condition (12). If the loss operator satisfies (6) (or equivalently (7)) mentioned in Definition 5, for some $\alpha\in(0,\infty)$, then, $\overline{\text{{gen}}}_{\rho}(\mathcal{N})\leq\sqrt{2\alpha^{2}\mathcal{I}_{bound}},$ where $\mathcal{I}_{bound}=\left[(|\mathcal{Z}|-1)\ln\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)\right]$ is the upper bound on the mutual information derived in Theorem 4.
@@ V-A Untrusted Data Processor Model
In this setting, the Respondent does not grant the Data Processor access to the raw classical data $s$ directly. Instead, the Respondent provides access only to the set of encoded quantum states $\{\rho_{s}\}_{s}$. The Processor is tasked with running a learning algorithm to produce a hypothesis $w$. Since the Data Processor is untrusted, the learning algorithm is modeled as a single, fixed completely positive, trace-preserving (CP-TP) map $\mathcal{N}$ that must be independent of the specific input index $s$. The total state generated at the output of this process is, $\mathcal{N}(\rho_{s}):=\sum_{w\in\mathcal{W}}\mathcal{N}_{w}(\rho_{s})\otimes|w\rangle\langle w|,$ where each $\mathcal{N}_{w}$ is a completely positive trace non-increasing map summing to $\mathcal{N}$. To ensure privacy, the Respondent mandates that this map $\mathcal{N}$ must satisfy differential privacy constraints with respect to neighboring inputs. The formal definition of such a $1$-neighbor $(\varepsilon,\delta)$-DP support-consistent learning algorithm for an untrusted Data Processor is provided in Definition 11 of Appendix G.
@@ V-B Information-Theoretic Admissibility (ITA): Motivation and Definition
If such a scenario is possible, the privacy guarantees calculated based on $\mathcal{N}$ (as mentioned in Definition 11) are rendered void, as the Data Processor effectively holds the information content of $\mathcal{N}^{\prime}$. To formalize this, we introduce the concept of ordering of informativeness between learning algorithms (Definition 12, Appendix G). We say that an algorithm $\mathcal{N}^{\prime}$ is more informative than $\mathcal{N}$ with respect to the ensemble $\{\rho_{s}\}_{s}$ if there exists a post-processing CP-TP map $\Gamma$ (a "simulation map") such that,
@@ V-C Quantum Advantage: Privacy under ITA
The Collapse of Classical Privacy: In the classical domain, the encoded states $\rho_{s}$ effectively behave as probability distributions (or commutative states). In such a scenario, the lemma below shows that simultaneous ITA (optimality) and Differential Privacy (indistinguishability) are impossible.
@@
See Appendix I
The implication of Lemma 1 is severe: classical ITA algorithms permit full reconstruction of the raw data. Since under ITA the Data Processor effectively holds the raw data, the output-based guarantees of Definition 11 are insufficient. Therefore, for classical ITA algorithms, the privacy condition in Definition 11 must be strengthened to require indistinguishability on the raw states $\rho_{s}$ directly, effectively substituting $\mathcal{N}(\rho_{s})$ with $\rho_{s}$. We discuss more on this in Appendix G-A.
Quantum Privacy via Non-Commutativity: In the quantum regime, the raw inputs $\{\rho_{s}\}_{s}$ may be non-commuting. Quantum mechanics dictates that non-orthogonal states cannot be perfectly distinguished, even by the optimal measurement. Therefore, an algorithm can be ITA—extracting the maximum physically permissible information—without yielding enough information to identify $s$ perfectly.
@@
Consider the states $\rho_{z}=|\phi_{z,p}\rangle\langle\phi_{z,p}|$ with $|\phi_{z,p}\rangle=\sqrt{1-p}$ $|0\rangle$ $(-1)^{z}\sqrt{p}|1\rangle$. For a dataset $s=(z_{1},\dots,z_{n})$, define the encoded state $|e_{s}\rangle:=\bigotimes_{j=1}^{n}|\phi_{z_{j},1/2}\rangle$. Let $P_{k}$ be the projection onto the subspace spanned by $\{|e_{s}\rangle:|s|_{1}=k\}$ (i.e., the subspace of states corresponds to all strings $s$ with Hamming weight $k$). The untrusted Data Processor applies the projective measurement map $\{\mathcal{N}_{k}\}_{k}$ defined by $\mathcal{N}_{k}(\rho):=P_{k}\rho P_{k}$.